# Efficient Feature Aggregation and Scale-Aware Regression for Monocular 3D Object Detection

### v2

@@ IV-D Visualization
In Fig. 9, we compare the proposed MonoASRH with the Transformer-based MonoDETR[48] and center-based MonoLSS[17]. Each group of images shows the camera view and BEV view. Compared to the other methods, MonoASRH produces higher-quality 3D bounding boxes across various scenes. Notably, it excels at detecting and localizing small distant objects, such as the white vehicle on the right in the third group of images, demonstrating the benefit of incorporating scale information. MonoASRH also provides more accurate results for fully visible objects. Additionally, it more effectively detects partially occluded objects, as shown on the left in the second group of images.

### v3

@@ IV-D Visualization
In Fig. 11, we compare the proposed MonoASRH with the Transformer-based MonoDETR[zhang2023monodetr] and center-based MonoLSS[li2024monolss]. Each group of images shows the camera view and BEV view. Compared to the other methods, MonoASRH produces higher-quality 3D bounding boxes across various scenes. Notably, it excels at detecting and localizing small distant objects, such as the white vehicle on the right in the third group of images, demonstrating the benefit of incorporating scale information. MonoASRH also provides more accurate results for fully visible objects. Additionally, it more effectively detects partially occluded objects, as shown on the left in the second group of images.